OpenStudy (anonymous):

The area of Australia/Oceania is approximately 7.6x10^6 square kilometers. Its population is approximately 3.11x10^7 people. What is the approximate population density (people per square kilometer) of Australia/Oceania? Write your answer in standard form. If necessary, round your answer to the nearest hundredth.

OpenStudy (campbell_st):

take the population and divide by the area \[Density = \frac{3.11 \times 10^7}{7.6 \times 10^6}\]

OpenStudy (anonymous):

ty